OpenStudy (anonymous):

Will medal right answer y=x-4 a( -1,-5) b(0,-3) c(2,6) d(3,1

OpenStudy (anonymous):

Hmm,what do u think? @meeep

OpenStudy (anonymous):

-1 -5

OpenStudy (anonymous):

correct!

OpenStudy (anonymous):

when u substitute x=-1 into the equation,u will get y=-5

OpenStudy (anonymous):

I have one more question if its ok

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

i will try my best

OpenStudy (anonymous):

which equation has (2,-1) as a solution a( y=2x-1 B(y=x+3 c(y=x-3 d(y=-2x - 1

OpenStudy (anonymous):

try substitute x=2 into each equations to find the value of y

OpenStudy (anonymous):

how I dont under stand how to do that

OpenStudy (anonymous):

given that the coordinates is (2,-1) which is (x,y)

OpenStudy (anonymous):